# Non-local Hamilton–Jacobi-Bellman equations for the stochastic optimal control of path-dependent piecewise deterministic processes

### v1

@@ 4.3. Path-dependent controlled PDPs on $\check{\Omega}$
for each $\check{\omega}=(t_{n},e_{n},\alpha_{n})_{n\in{\mathbb{N}}_{0}}\in\check{\Omega}$. Keep in mind that, by Remark 2.1, the $X^{s,x}$-term on the right-hand side of (4.2) depends only on its values on $[0,t_{n}]$, i.e., our definition of $X^{s,x}$ is not circular. Note that ${X}^{s,x}(\check{\omega})\in\Omega$ for each $\check{\omega}\in\check{\Omega}$ with $t_{0}<\infty$. Also note that in (4.2), $t_{n+1}=\infty$ for some $n\in{\mathbb{N}}_{0}$ is possible and if, in addition, $t_{n}<\infty$, then $X^{s,x}(t,\check{\omega})=\phi^{t_{n},X^{s,x}(\check{\omega})\otimes_{t_{n}}\,e_{n},\alpha_{n}}(t)$ for all $t\in[t_{n},\infty)$. In particular, $X^{s,x}(t,(t_{0},e_{0},\alpha_{0};\infty,\Delta,\Delta^{\prime};\infty,\Delta,\Delta^{\prime};\ldots))=\phi^{s,x}(t)$ if $t_{0}<\infty$.
@@
We also consider the following “restrictions” of $X^{s,x}$. Given $n\in{\mathbb{N}}_{0}$, define $X^{s,x}_{n}:{\mathbb{R}}_{+}\times\check{\Omega}_{n+1}\to E\cup\{\Delta\}$ by
@@
Denote by $\mathbb{A}^{\prime}$ the space of all policies $\mathbf{a}=(\mathbf{a}_{0},\mathbf{a}_{1},\ldots)$ such that each $\mathbf{a}_{n}=\mathbf{a}_{n}({\mathrm{d}}\alpha_{n}|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n-1})$, $n\in{\mathbb{N}}_{0}$, is a universally measurable121212 Recall from Definition 7.18 in [BS] that a subset $\underline{S}$ of a Borel space $S$ is universally measurable if, for every ${{\mathbb{P}}}\in\mathscr{P}(S)$, the set $\underline{S}$ belongs to the completion of $\mathcal{B}(S)$ with respect to ${{\mathbb{P}}}$. stochastic kernel on $\mathcal{A}\cup\{\Delta^{\prime}\}$ given ((+×$(({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\})^{n+1}\times(\mathcal{A}\cup\{\Delta^{\prime}\})^{n}$.
Denote by $\mathbb{A}$ the space of all non-randomized policies $\mathbf{a}=(\mathbf{a}_{0},\mathbf{a}_{1},\ldots)\in\mathbb{A}^{\prime}$, i.e., for each $n\in{\mathbb{N}}_{0}$, $(t_{0},e_{0})$, $\ldots$, $(t_{n},e_{n})\in({\mathbb{R}}_{+}\times E)\cup\{\infty,\Delta\}$, and $\alpha_{0}$, $\ldots$, $\alpha_{n-1}\in\mathcal{A}\cup\{\Delta^{\prime}\}$, $\mathbf{a}_{n}({\mathrm{d}}\alpha_{n}|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n-1})$ is a Dirac measure.
@@
We want to stress that our use of randomized policies is mainly done for the sake of simplicity. While the setting appears to become more complex compared to the use of path-dependent feedback controls, the notation actually becomes easier and the embedding of our continuous-time model into a discrete-time model becomes straight-forward (see also Remark A.2).
@@
Let $(s,x)\in(0,T)\times\Omega$ and $\mathbf{a}\in\mathbb{A}^{\prime}$. Construct a probability measure ˇ$\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}$ on $(\check{\Omega},\check{{\mathcal{F}}})$ with corresponding expected value $\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}$ recursively by defining its marginals ˇ$\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}_{n}$ on $(\check{\Omega}_{n},\check{{\mathcal{F}}}_{n,\infty})$ with corresponding expected values $\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{n}$, $n\in{\mathbb{N}}$, as follows (cf. section 3.2 of [Jacobsen]):
@@
where $\psi_{n}:({\mathbb{R}}_{+}\times E)^{n+1}\times\mathcal{A}^{n+1}\to{\mathbb{R}}$, $n\in{\mathbb{N}}_{0}$, are Borel-measurable and bounded and the notation $\check{\omega}_{n+1}=(t_{j},e_{j},\alpha_{j})_{j=0}^{n}$ is used.
4.4. The optimal control problem
@@
The proofs of the next two theorems are in section 4.5.

### v2

@@ 4.3. The continuous-time optimal control problem
for each $\check{\omega}=(t_{n},e_{n},\alpha_{n})_{n\in{\mathbb{N}}_{0}}\in\check{\Omega}$.
Remark 4.7.
Keep in mind that, by Remark 2.1, the $X^{s,x}$-term on the right-hand side of (4.3) depends only on its values on $0,t_{n}]$, i.e., our definition of $X^{s,x}$ is not circular. Note that ${X}^{s,x}(\check{\omega})\in\Omega$ for each $\check{\omega}\in\check{\Omega}$ with $t_{0}<\infty$. Also note that in (4.3), $t_{n+1}=\infty$ for some $n\in{\mathbb{N}}_{0}$ is possible and if, in addition, $t_{n}<\infty$, then $X^{s,x}(t,\check{\omega})=\phi^{t_{n},X^{s,x}(\check{\omega})\otimes_{t_{n}}\,e_{n},\alpha_{n}}(t)$ for all $t\in[t_{n},\infty)$. In particular, $X^{s,x}(t,(t_{0},e_{0},\alpha_{0};\infty,\Delta,\Delta^{\prime};\infty,\Delta,\Delta^{\prime};\ldots))=\phi^{s,x}(t)$ if $t_{0}<\infty$.
@@
We also consider the following “restrictions” of $X^{s,x}$. Given $n\in{\mathbb{N}}_{0}$, define $X^{s,x}_{n}:{\mathbb{R}}_{+}\times\check{\Omega}_{n+1}\to{\mathbb{R}}^{d}\cup\{\Delta\}$ by
@@
Next, we introduce classes of policies needed to formulate our optimal control problem. Note that we consider randomized as well as non-randomized policies. We use randomized policies mainly for technical reasons: Doing so will allow us to directly apply results from discrete-time stochastic optimal control in [BS]. We will show that for our problem optimization over randomized policies and optimization over non-randomized policies are equivalent.
@@
Denote by $\mathbb{A}^{\prime}$ the space of all policies $\mathbf{a}=(\mathbf{a}_{0},\mathbf{a}_{1},\ldots)$ such that each $\mathbf{a}_{n}=\mathbf{a}_{n}({\mathrm{d}}\alpha_{n}|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n-1})$, $n\in{\mathbb{N}}_{0}$, is a universally measurable stochastic kernel on $\mathcal{A}\cup\{\Delta^{\prime}\}$ given $(({\mathbb{R}}_{+}\times{\mathbb{R}}^{d})\cup\{(\infty,\Delta)\})^{n+1}\times(\mathcal{A}\cup\{\Delta^{\prime}\})^{n}$.
Denote by $\mathbb{A}$ the space of all non-randomized policies $\mathbf{a}=(\mathbf{a}_{0},\mathbf{a}_{1},\ldots)\in\mathbb{A}^{\prime}$, i.e., for each $n\in{\mathbb{N}}_{0}$, $(t_{0},e_{0})$, $\ldots$, $(t_{n},e_{n})\in({\mathbb{R}}_{+}\times{\mathbb{R}}^{d})\cup\{\infty,\Delta\}$, and $\alpha_{0}$, $\ldots$, $\alpha_{n-1}\in\mathcal{A}\cup\{\Delta^{\prime}\}$, $\mathbf{a}_{n}({\mathrm{d}}\alpha_{n}|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n-1})$ is a Dirac measure.
@@
Given an initial condition and a specified policy, we are now in the position to define the corresponding probability measure.
@@
Let $(s,x)\in(0,T)\times\Omega$ and $\mathbf{a}\in\mathbb{A}^{\prime}$. We construct a probability measure $\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}$ on $(\check{\Omega},\check{{\mathcal{F}}})$ with corresponding expected value $\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}$ recursively by defining its marginals $\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}_{n}$ on $(\check{\Omega}_{n},\check{{\mathcal{F}}}_{n,\infty})$ with corresponding expected values $\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{n}$, $n\in{\mathbb{N}}$, as follows (cf. section 3.2 of [Jacobsen]):
@@
where $\psi_{n}:({\mathbb{R}}_{+}\times{\mathbb{R}}^{d})^{n+1}\times\mathcal{A}^{n+1}\to{\mathbb{R}}$, $n\in{\mathbb{N}}_{0}$, are Borel-measurable and bounded and the notation $\check{\omega}_{n+1}=(t_{j},e_{j},\alpha_{j})_{j=0}^{n}$ is used.
Thanks to the Kolmogorov extension theorem, the above construction really provides a unique probability measure (besides section 3.2 in [Jacobsen], we refer also to Proposition 7.45 and its proof in [BS] for a detailed treatment).
Definition 4.12.
@@
4.4. A related discrete-time optimal control problem and the dynamic programming principle
We will construct a path-dependent discrete-time infinite-horizon decision model, which is closely related to our continuous-time problem. This will enable us to establish a dynamic programming principle for the value functions $V$ and $V^{\prime}$ of our continuous-time problem (see Definition 4.12). (Our approach is inspired by similar work in the Markovian case, see, e.g., section 3 of [BR10PDMP], section 8.2 in [BR11book], and [RTT17PDMP].)
In this subsection, we fix $(s,x)\in(0,T]\times\Omega$.
We consider a path-dependent infinite-horizon model with the following data (cf. Definitions 8.1 and 9.1 in [BS] for the non-path-dependent case):
  • State space: $S=({\mathbb{R}}_{+}\times{\mathbb{R}}^{d})\cup\{(\infty,\Delta)\}$.
  • Control space: $U=\mathcal{A}\cup\{\Delta^{\prime}\}$.
  • Disturbance space: $S$.
  • Disturbance kernel
    $\vec{Q}_{n}({\mathrm{d}}t\,{\mathrm{d}}e|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n})\text{ at stage $n$, $n\in{\mathbb{N}}_{0}$,}$
    is a Borel-measurable stochastic kernel on $S$ given $S^{n+1}\times U^{n+1}$ defined by
    $\int_{S}{\mathbf{1}}_{[0,\infty)}(t)\,\psi(t,e)\,\vec{Q}_{n}({\mathrm{d}}t\,{\mathrm{d}}e|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n})$
    $:=\begin{cases}&-\int_{{\mathbb{R}}_{+}\times E}\psi(t,e)\,Q(t,{\mathrm{d}}e|t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n})\,{\mathrm{d}}F(t|t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n})\\ &\phantom{\psi(\infty,\Delta)}\qquad\text{if $\check{\omega}_{n+1}=(t_{j},e_{j},\alpha_{j})_{j=0}^{n}\in\check{\Omega}_{n+1}$,}\\ &\psi(\infty,\Delta)\qquad\text{otherwise,}\end{cases}$
    for every bounded Borel-measurable function $\psi:S\to{\mathbb{R}}$.
  • Cost function
    $g_{n}:S^{n+1}\times U^{n+1}\to{\mathbb{R}}_{+}\cup\{\infty\}\quad\textup{at stage $n$, $n\in{\mathbb{N}}_{0}$,}$
    is a lower semi-analytic function defined by
    $g_{n}((t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n}):=\begin{cases}\vec{\ell}_{n}((t_{j},e_{j},\alpha_{j})_{j=0}^{n})\text{ if $(t_{j},e_{j},\alpha_{j})_{j=0}^{n}\in\check{\Omega}_{n+1}$,}\\ \infty\text{ otherwise,}\end{cases}$
    where $\vec{\ell}_{n}:\check{\Omega}_{n+1}\to{\mathbb{R}}_{+}$ is defined by
Remark 4.13.
Borel-measurability of $\vec{Q}_{n}$ and lower semi-analyticity of $g_{n}$, $n\in{\mathbb{N}}_{0}$, follow from Remarks 4.3, 4.6, and 4.8. Keep in mind that $(s,x)$ is fixed here.
Depending on the starting stage $k\in{\mathbb{N}}_{0}$, we use as policy space the set $\mathbb{A}^{k}$ of all policies $\mathbf{a}^{k}=(\mathbf{a}^{k}_{k},\mathbf{a}^{k}_{k+1},\ldots)$ such that each
$\mathbf{a}^{k}_{j}=\mathbf{a}^{k}_{j}({\mathrm{d}}\alpha_{j}|(t_{0},e_{0}),\ldots,(t_{j},e_{j});\alpha_{0},\ldots,\alpha_{j-1}),\quad j\in\{k,k+1,\ldots\},$
is a universally measurable stochastic kernel on $U$ given $S^{j+1}\times U^{j}$ ([BS, section 10.1]).
Remark 4.14.
Note that $\mathbb{A}^{0}$ coincides with $\mathbb{A}^{\prime}$ from Definition 4.9.
Given $k\in{\mathbb{N}}_{0}$, a policy $\mathbf{a}^{k}=(\mathbf{a}^{k}_{k},\mathbf{a}^{k}_{k+1},\ldots)\in\mathbb{A}^{k}$, and a probability measure $p^{k}$ on $S^{k+1}\times U^{k}$, define marginals $r^{k}_{N}(\mathbf{a}^{k},p^{k})$ on $(S\times U)^{N}$, $N\in\{k+1,k+2,\ldots\}$ by
(4.8) $\begin{split}&\int\psi\,{\mathrm{d}}r^{k}_{N}(\mathbf{a}^{k},p^{k})\\ &:=\int_{S^{k+1}\times U^{k}}\int_{U}\int_{S}\cdots\int_{S}\int_{U}\psi((t_{0},e_{0}),\ldots,(t_{N-1},e_{N-1});\alpha_{0},\ldots,\alpha_{N-1})\\ &\qquad\qquad\mathbf{a}^{k}_{N-1}({\mathrm{d}}\alpha_{N-1}|(t_{0},e_{0}),\ldots,(t_{N-1},e_{N-1});\alpha_{0},\ldots,\alpha_{N-2}))\\ &\qquad\qquad\vec{Q}_{N-1}({\mathrm{d}}t_{N-1}\,{\mathrm{d}}e_{N-1}|(t_{0},e_{0}),\ldots,(t_{N-2},e_{N-2});\alpha_{0},\ldots,\alpha_{N-2})\cdots\\ &\qquad\qquad\vec{Q}_{k}({\mathrm{d}}t_{k+1}\,{\mathrm{d}}e_{k+1}|(t_{0},e_{0}),\ldots,(t_{k},e_{k});\alpha_{0},\ldots,\alpha_{k})\\ &\qquad\qquad\mathbf{a}^{k}_{k}({\mathrm{d}}\alpha_{k}|(t_{0},e_{0}),\ldots,(t_{k},e_{k});\alpha_{0},\ldots,\alpha_{k-1})\\ &\qquad\qquad p^{k}({\mathrm{d}}t_{0}\,{\mathrm{d}}e_{0}\ldots{\mathrm{d}}t_{k}\,{\mathrm{d}}e_{k}\,{\mathrm{d}}\alpha_{0}\ldots{\mathrm{d}}\alpha_{k-1})\end{split}$
for each bounded universally measurable function $\psi:S^{N}\times U^{N}\to{\mathbb{R}}$.
For each $k\in{\mathbb{N}}_{0}$, $\mathbf{a}^{k}\in\mathbb{A}^{k}$, $(t_{0},e_{0})$, $\ldots$, $(t_{k},e_{k})\in S$, and $\alpha_{0}$, $\ldots$, $\alpha_{k-1}\in U$, define the corresponding cost
(4.9) $\begin{split}&J_{\mathbf{a}^{k}}(k;(t_{0},e_{0}),\ldots,(t_{k},e_{k});\alpha_{0},\ldots,\alpha_{k-1})\\ &:=\sum_{j=k}^{\infty}\int g_{j}\,{\mathrm{d}}r_{j+1}^{k}(\mathbf{a}^{k},\delta_{((t_{0},e_{0}),\ldots,(t_{k},e_{k});\alpha_{0},\ldots,\alpha_{k-1})}).\end{split}$
For each $k\in{\mathbb{N}}_{0}$, $(t_{0},e_{0})$, $\ldots$, $(t_{k},e_{k})\in S$, and $\alpha_{0}$, $\ldots$, $\alpha_{k-1}\in U$, define the corresponding optimal cost
(4.10) $\begin{split}&J^{\ast}(k;(t_{0},e_{0}),\ldots,(t_{k},e_{k});\alpha_{0},\ldots,\alpha_{k-1})\\ &:=\inf_{\mathbf{a}^{k}\in\mathbb{A}^{k}}J_{\mathbf{a}^{k}}(k;(t_{0},e_{0}),\ldots,(t_{k},e_{k});\alpha_{0},\ldots,\alpha_{k-1}).\end{split}$
Next, we show that, for appropriate values, the just defined optimal cost function $J^{\ast}$ coincides with the value function $V^{\prime}$ defined by (4.6).
Theorem 4.15.
Fix $(s,x)\in(0,T)\times\Omega$ and $(t,e,\alpha)\in(s,T]\times{\mathbb{R}}^{d}\times\mathcal{A}$. Then
(4.11) $V^{\prime}(s,x)$ $=J^{\ast}(0;s,x(s)),$
(4.12) $V^{\prime}(t,\phi^{s,x,\alpha}\otimes_{t}e)$ $=J^{\ast}(1;(s,x(s)),(t,e);\alpha).$
Proof.
(i) First, we establish (4.11). To this end, let $\mathbf{a}\in\mathbb{A}^{\prime}$. Then, by (4.6),
$J(s,x,\mathbf{a})$ $=\sum_{n=0}^{\infty}\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}\Bigl[\int_{T_{n}}^{T_{n+1}}{\mathbf{1}}_{[0,T]}(\tau)\,\ell(\tau,X^{s,x},\mathcal{A}_{n}(\tau))\,{\mathrm{d}}\tau+{\mathbf{1}}_{[T_{n},T_{n+1})}(T)\,h(X^{s,x})\Bigr].$
${\color[rgb]{0,0,0}{\mathbf{1}}_{(T,\infty]}(T_{n+1})\,h(X^{s,x}(\check{\omega}))={\mathbf{1}}_{(T,\infty]}(T_{n+1})\,h(X^{s,x}_{n}(\check{\omega}_{n+1}))}$
$\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}\Bigl[\int_{T_{n}}^{T_{n+1}}{\mathbf{1}}_{[0,T]}(\tau)\,\ell(\tau,X^{s,x},\mathcal{A}_{n}(\tau))\,{\mathrm{d}}\tau+{\mathbf{1}}_{[T_{n},T_{n+1})}(T)\,h(X^{s,x})\Bigr]$
$=\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{n+2}\Bigl[\int_{T_{n}\wedge T}^{T}{\mathbf{1}}_{(\tau,\infty]}(T_{n+1})\ell(\tau,X_{n}^{s,x},\mathcal{A}_{n}(\tau))\,{\mathrm{d}}\tau+{\mathbf{1}}_{[0,T]}(T_{n}){\mathbf{1}}_{(T,\infty]}(T_{n+1})\,h(X_{n}^{s,x})\Bigr]$
$=\int_{\check{\Omega}_{n+1}}\Big(\int_{t_{n}\wedge T}^{T}\chi^{t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(\tau)\,\ell(\tau,X^{s,x}_{n}(\check{\omega}_{n+1}),a_{n}(\tau))\,{\mathrm{d}}\tau$
Thus, recalling the definition of $\vec{\ell}_{n}$ in (4.7), we have
(4.13) $J(s,x,\mathbf{a})=\sum_{n=0}^{\infty}\check{{{\mathbb{E}}}}_{n+1}^{s,x,\mathbf{a}}\left[\vec{\ell}_{n}\right].$
By the definitions of the marginals $r^{k}_{N}(\mathbf{a}^{k},p)$ in (4.8) and $\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}_{n}$ in (4.5),
$\int g_{0}\,{\mathrm{d}}r^{0}_{1}\left(\mathbf{a},\delta_{(s,x(s)}\right)$ $=\int_{S\setminus(\{0\}\times E)}\int_{U}\vec{\ell}_{0}(t_{0},e_{0},\alpha_{0})\,\mathbf{a}_{0}({\mathrm{d}}\alpha_{0})\,\delta_{(s,x(s))}({\mathrm{d}}t_{0}\,{\mathrm{d}}e_{0})$
$\qquad+\infty.\delta_{(s,x(s))}(\{0\}\times E)\qquad\text{(note that $s>0$)}$
$=\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{1}\left[\vec{\ell}_{0}\right],$
$\int g_{1}\,{\mathrm{d}}r^{0}_{2}\left(\mathbf{a},\delta_{(s,x(s))}\right)$ $=\int_{S}\int_{U}\int_{S}\int_{U}g_{1}((t_{0},e_{0}),(t_{1},e_{1});\alpha_{0},\alpha_{1})$
$=\check{{{\mathbb{E}}}}^{s,\mathbf{a}}_{2}\left[\vec{\ell}_{1}\right]\qquad\text{because $\left[r^{0}_{2}\left(\mathbf{a},\delta_{(s,x(s))}\right)\right](\check{\Omega}_{2})=1$,}$
$\int g_{N}\,{\mathrm{d}}r^{0}_{N+1}\left(\mathbf{a},\delta_{(s,x(s))}\right)$ $=\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{N+1}\left[\vec{\ell}_{N}\right]\qquad\text{(can be shown by induction)}.$
Hence, by (4.13) and (4.9), we have
(4.14) $J(s,x,\mathbf{a})=J_{\mathbf{a}}(0;s,x(s)).$
Since $\mathbf{a}$ was an arbitrary policy in $\mathbb{A}^{\prime}=\mathbb{A}^{0}$, we can conclude that (4.11) holds.
(ii) Now, we prove (4.12). Let
$y:=\phi^{s,x,\alpha}\otimes_{t}e$
(4.15) $J(t,y,\mathbf{a})=\sum_{n=0}^{\infty}\check{{{\mathbb{E}}}}^{t,y,\mathbf{a}}_{n+1}\left[\vec{\ell^{1}_{n}}\right],$
(4.16) $\begin{split}\vec{\ell_{n}^{1}}(\check{\omega}_{n+1})&:=\int_{t_{n}\wedge T}^{T}\chi^{t_{n},X^{t,y}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(\tau)\,\ell(t,X^{t,y}_{n}(\check{\omega}_{n+1}),\alpha_{n}(\tau))\,{\mathrm{d}}\tau\\ &\qquad+{\mathbf{1}}_{[0,T]}(t_{n})\,\chi^{t_{n},X^{t,y}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(T)\,h(X^{t,y}_{n}(\check{\omega}_{n+1}))\end{split}$
First note that, for every $\mathbf{a}^{1}\in\mathbb{A}^{1}$,
$=\int_{U}\vec{\ell}_{1}((s,x(s),\alpha),(t,e,\alpha_{1}))$
[…]
Together with the definition of $\vec{\ell_{0}^{1}}$ in (4.16), we obtain
[…]
Next, let $N\in\{2,3,\ldots\}$ and $\mathbf{a}^{1}\in\mathbb{A}^{1}$. Then
(4.17) $\begin{split}&g_{N}((s,x(s)),(t,e),(t_{2},e_{2}),\ldots,(t_{N},e_{N});\alpha,\alpha_{1},\ldots,\alpha_{N})\\ &\qquad=\vec{\ell}_{N}((s,x(s),\alpha),(t,e,\alpha_{1}),(t_{2},e_{2},\alpha_{2}),\ldots,(t_{N},e_{N},\alpha_{N}))\end{split}$
and the right-hand side of (4.17) depends on
which equals
Thus
$\,\,\,\,\mathbf{a}_{1}^{1}({\mathrm{d}}\alpha_{1}|(s,x(s)),(t,y(t));\alpha)$
$=^{(\text{after relabeling of variables})}$
$\,\,\,\,\mathbf{a}_{1}^{1}({\mathrm{d}}\alpha_{0}|(s,x(s)),(t,y(t));\alpha)$
Consequently, we have
[…]
This concludes the proof. ∎
Theorem 4.16.
Let $(t_{0},e_{0})\in S$. Then
$J^{\ast}(0;t_{0},e_{0})=\inf_{\alpha\in U}\Bigl\{g_{0}((t_{0},e_{0});\alpha)+\int_{S}J^{\ast}(1;(t_{0},e_{0}),(t_{1},e_{1});\alpha)\,\vec{Q}_{0}({\mathrm{d}}t_{1}\,{\mathrm{d}}e_{1}|(t_{0},e_{0});\alpha)\Bigr\}.$
Proof.
This result is a specific instance of Theorem A.3, whose proof is provided in appendix A. ∎